# SimSort: A Data-Driven Framework for Spike Sorting by Large-Scale Electrophysiology Simulation

### v2

@@ 1 Introduction
Understanding the complex computations performed by the brain requires insight into the activity of individual neurons [1, 2], which is crucial for exploring how information is encoded, processed, and transmitted within neural circuits, as well as decoding the brain’s functional dynamics [3, 4, 5]. Recent advances in neural recording technologies have enabled capturing the activity of neurons across multiple regions of the brain with high spatial and temporal precision [6, 7, 8, 9]. Notably, extracting meaningful information from recordings relies on a critical step known as spike sorting.
Spike sorting is the process of extracting and identifying neural activity from extracellular recordings. It involves two main steps (Fig. 1): spike detection, which extracts spike events from background noise, and spike identification, which assigns these detected spikes to individual neurons [10]. Spike sorting is indispensable for transforming raw electrical signals into interpretable data that reveal the firing patterns of individual neurons. Accurate spike sorting is essential for linking neural activity to behavior, understanding the functional organization of neural circuits [11], and uncovering mechanisms underlying various sensory and cognitive processes [12]. Furthermore, its role extends to translational applications, such as improving neural prosthetics and developing closed-loop BCIs, where spike sorting is critical for achieving precise neural decoding and control [13, 14].
For a long time, spike sorting predominantly relied on heuristic statistics and machine learning approaches [1, 15, 16, 17, 6, 18, 19]. While these methods have promoted neuroscience research, they exhibit several key limitations. First, their sorting results are sensitive to parameter settings and post-processing, which depend on the experimenter’s expertise and must be customized for each dataset. Moreover, these approaches lack a data-driven foundation, limiting their scalability and adaptability across diverse experimental settings, particularly in low signal-to-noise ratio (SNR) and high-variability scenarios, where their performance declines significantly. Recently, deep learning-based methods, such as YASS [20] and CEED [21], have attempted to improve spike sorting by adopting data-driven approaches to enhance detection and clustering accuracy. However, their generalizability and practical use are hindered by the limited training datasets.
@@ 2 Preliminaries
The goal of spike sorting (Figure 1) is to extract single-neuron spiking activity from extracellular recordings, where the activities of an unknown number of neurons are mixed together. Mathematically, spike sorting is a blind source separation problem [22] and can be formally defined as follows: Let $\mathbf{V}\in\mathbb{R}^{T\times C}$ represent the extracellular recording (voltage) across $C$ electrode channels over $T$ time points. The goal is to infer a set of spike times $\{t_{k}\}_{k=1}^{K}$ and corresponding neuron labels $\{y_{k}\}_{k=1}^{K}$, where $y_{k}\in\{1,\dots,N\}$ represents the neuron identity, and $N$ is the total number of detected neurons.
@@ 3 Related Work
Early spike sorting pipelines typically used thresholds to detect spike events, followed by dimensionality reduction (e.g., PCA or $t$-SNE) and clustering [23, 24, 25]. Although this pipeline was straightforward, it often suffered from noise susceptibility, inaccurate detection of low SNR events, and reliance on manual parameter tuning.
Deep learning approaches have been adopted to replace or enhance certain steps in traditional pipelines, aiming to improve robustness, automation, and adaptability to diverse experimental conditions. Autoencoders have been explored as a method for dimensionality reduction [26, 27, 28]. YASS [20] employed a convolutional neural network for spike detection and waveform cleaning, thereby mitigating clustering errors caused by distorted waveforms. CEED [21] applied a contrastive learning framework to enforce invariances to amplitude fluctuations, noise, and channel subset changes in the extracted embeddings. However, these methods were limited by their reliance on restricted training datasets, potentially reducing their generalization across diverse experimental conditions.
@@ 4.1 Dataset Creation
To address the scarcity of labeled spike data, we generated a large-scale synthetic dataset using biophysically detailed neuron models from the Blue Brain Project (BBP) [29], covering 206 neuron models from layers 1–6 of the juvenile rat somatosensory cortex. These models capture diverse morphologies and electrophysiological dynamics across 30 neuronal types.
@@
Based on these models, we first simulated intracellular activity by injecting noise currents into multi-compartment neurons using the NEURON simulator [30], producing realistic spiking responses. Then, extracellular potentials were computed via volume conductor theory, summing transmembrane currents across compartments to obtain virtual recordings from tetrode electrodes randomly placed near the neurons (Figure 2; full technical details in Appendix A).
Each simulation trial included five neurons of varying types placed in a $100\times 100\times 100~{}\mu$m3 volume, with randomized positions for both cells and electrodes. We generated 8,192 trials in total, yielding over 40,000 annotated neuron instances. Ground-truth spike times were extracted from intracellular traces and aligned to extracellular recordings.
@@
We generated a large-scale dataset consisting of 8192 simulated recording trials, representing continuous neuronal activities from more than 40,000 individual neurons. Each trial was configured with randomized neurons and electrode positions, along with diverse combinations of neuron types, to ensure high variability and diversity. Intracellular spike timestamps were recorded as ground-truth for extracellular spikes, providing precise annotations for model training and evaluation.
@@ 4.2 Spike Detection
We utilized a Transformer-based architecture [31] for spike detection, trained on the simulated continuous signal dataset detailed in Sec. 4.1. For robustness and generalization, data augmentation was integrated into the training process, referring to Fig. 3. We employed a binary cross-entropy loss for the spike detection model:
@@ 4.3 Spike Identification
Contrastive Learning: We leveraged contrastive learning to learn representations that capture the relationships between spike waveforms (Fig. 4). Contrastive learning sought to map neural data into an embedding space where related examples were positioned nearby, while unrelated examples were placed further apart [32, 33]. For each waveform $X_{i}$, we constructed a triplet by selecting $X_{i}^{+}$ with the same label and $X_{i}^{-}$ with a different label. The objective was [34]:
@@
Denoising: The augmented waveform $X_{i}^{\prime}\in\mathbb{R}^{L\times C}$ then underwent a denoising step using Singular Value Decomposition (SVD) [35, 36]. The waveform was first reshaped and decomposed as $X_{i}^{\prime}=U\Sigma V^{\top}$, where $U$, $\Sigma$, and $V$ represented the singular vectors and singular values. The first $k$ components were retained to reconstruct the waveform.
Encoder: The denoised waveform $X_{i,\text{recon}}^{\prime}$ was fed into a Gated Recurrent Unit (GRU) encoder [37] (Fig. 4), which processed the input sequentially. The final hidden state $h_{T}$ served as the learned representation $f(X_{i})$, encapsulating the spatiotemporal features of the waveform for subsequent clustering tasks.
Clustering: After obtaining the waveform representation, we applied dimensionality reduction using UMAP [38] before clustering (Fig. 4). We tested two clustering approaches: a parametric algorithm, Gaussian Mixture Model (GMM) [39], and a non-parametric one, Mean Shift (MS) [40].
@@ 5.1 Benchmarks
  • Real-world data-based synthetic datasets (with label): Hybrid [41], WaveClus [42], IBL Neuropixels [21]. We evaluate both zero-shot (using only our simulated dataset) and fine-tuned (using also the dataset) performance of SimSort.
@@
Evaluation Metrics. To evaluate the performance of the overall spike sorting and spike detection tasks, we used Accuracy, Recall, and Precision as the evaluation metrics. Note that these metrics are specifically defined for the spike sorting context, where they measure the alignment between detected spikes and ground truth spike events, rather than their conventional definitions in classification tasks. For the spike identification task, we used the Adjusted Rand Index (ARI) [43] to measure clustering accuracy relative to the ground truth.
Table 1: Spike sorting results on BBP L6 dataset. The results for other methods were obtained through SpikeInterface [44]. Values are as mean $\pm$ S.E.M.
@@ 5.2 Spike Sorting Results on simulated dataset
We first examined Simsort on a simulated test set (BBP L6). In Table 1, we evaluated SimSort alongside widely used algorithms, including KiloSort [17], KiloSort2, MountainSort4 [6], and MountainSort5, implemented via the SpikeInterface pipeline [44], on the simulated BBP L6 dataset. SimSort achieved the best performance across all metrics, demonstrating its robustness and strong zero-shot generalization ability under idealized, simulated conditions.
Table 2: Spike sorting results on Hybrid dataset. Data for other methods were sourced from SpikeForest [41]. Values are mean $\pm$ S.E.M. Best results are in bold. Note that SimSort did not use this dataset for training, indicating a zero-shot generalization.
Methods Hybrid-static (SNR$>$3, 9 recordings) Hybrid-drift (SNR$>$3, 9 recordings)
@@
SimSort 0.62 $\pm$ 0.04 0.68 $\pm$ 0.04 0.77 $\pm$ 0.03 0.56 $\pm$ 0.03 0.63 $\pm$ 0.03 0.69 $\pm$ 0.03
Table 3: Spike sorting results on WaveClus dataset. The results for other methods were obtained through SpikeInterface [44]. Values are mean $\pm$ S.E.M. Note that SimSort did not use this dataset for training, indicating a zero-shot generalization.
@@ 5.3 Spiking Sorting results on real-world data-based synthetic datasets
Given SimSort’s strong performance on simulated data, we extended our evaluation to real-world data-based datasets to further test its zero-shot generalizability and robustness (RQ1). On the Hybrid dataset, we compared SimSort against several algorithms via SpikeForest, including HerdingSpikes2 [18], IronClust, JRClust, KiloSort [17], KiloSort2, MountainSort4 [6], SpykingCircus, and Tridesclous. The results are summarized in Table 2 and Fig. S1. On this dataset, SimSort demonstrated the highest Accuracy and Precision in static recordings and achieved the best Accuracy and Recall under drift conditions, although it slightly trailed KiloSort in Precision.
@@ 5.5 Detailed Analysis
Threshold detector 0.61$\pm$0.03 0.71$\pm$0.02 0.81$\pm$0.02
Static SimSort’s detector 0.72$\pm$0.03 0.84$\pm$0.02 0.82$\pm$0.02
Threshold detector 0.60$\pm$0.03 0.70$\pm$0.03 0.80$\pm$0.02
Drift SimSort’s detector 0.68$\pm$0.03 0.82$\pm$0.02 0.81$\pm$0.02
@@
After detecting spikes, our spike identification model extracted waveform representations in the latent space, enabling subsequent clustering. To effectively demonstrate the superiority of our identification model, we assembled a waveform dataset by gathering waveform data from each neuron in every hybrid recording. Using this dataset, we compared the 2D UMAP embedding of our model’s inferred representations with features extracted directly from waveforms using widely adopted dimensionality reduction techniques, including PCA [23], $\boldsymbol{t}$-SNE [24], and UMAP [38]. For evaluation, we applied a parametric clustering method, GMM [39], to compute the Adjusted Rand Index (ARI), providing a quantitative measure of the reliability of the representations.
@@
Additionally, we compared SimSort with CEED [21], a recent framework for representation learning on extracellular data, using the author-provided model. Note that CEED is trained on datasets derived from IBL DanLab DY016 and DY009 Neuropixels recordings, which are more closely related to our test sets. We evaluated both SimSort and CEED on fifty test sets from the IBL Neuropixels dataset (specifically derived from the IBL CortexLab KS046 recording). As presented in Table S4, Fig. S2S3, our method, without training on the dataset, achieved performance comparable to CEED.
@@
Impact of Dimensionality Reduction and Clustering Methods: To evaluate the adaptability of SimSort, we conducted a study to assess the impact of different dimensionality reduction techniques (PCA, $t$-SNE, UMAP, and None) and clustering algorithms (GMM, KMeans [45], Mean Shift [40], and Spectral Clustering [46]). Dimensionality reduction was applied to latent features extracted from waveforms, with an additional setup where clustering was performed directly on the raw latent features without dimensionality reduction (None). As illustrated in Fig. S4, the combination of UMAP and Mean Shift achieved the highest accuracy on both hybrid static and drift subsets. Across many other combinations, SimSort also maintained competitive performance, demonstrating its flexibility and robustness to various configurations.
@@ 6 Conclusion
This work was supported partially by AI for Science Foundation of Fudan University (FudanX24AI046 to Y.G.). This work was also supported by Microsoft Research.

### v3

@@ 1 Introduction
Understanding the complex computations performed by the brain requires insight into the activity of individual neurons lewickiReviewMethodsSpike1998; buzsakiLargescaleRecordingNeuronal2004, which is crucial for exploring how information is encoded, processed, and transmitted within neural circuits, as well as decoding the brain’s functional dynamics quianquirogaExtractingInformationNeuronal2009; joshiDynamicSynchronizationHippocampal2023; sarkarAdvancedSpikeSorting2024. Recent advances in neural recording technologies have enabled capturing the activity of neurons across multiple regions of the brain with high spatial and temporal precision chungFullyAutomatedApproach2017; steinmetzNeuropixels20Miniaturized2021a; chungHighDensityLongLastingMultiregion2019; hongNovelElectrodeTechnologies2019. Notably, extracting meaningful information from recordings relies on a critical step known as spike sorting.
Spike sorting is the process of extracting and identifying neural activity from extracellular recordings. It involves two main steps (Fig. 1): spike detection, which extracts spike events from background noise, and spike identification, which assigns these detected spikes to individual neurons lefebvreRecentProgressMultielectrode2016. Spike sorting is indispensable for transforming raw electrical signals into interpretable data that reveal the firing patterns of individual neurons. Accurate spike sorting is essential for linking neural activity to behavior, understanding the functional organization of neural circuits sibilleHighdensityElectrodeRecordings2022, and uncovering mechanisms underlying various sensory and cognitive processes liuDecodingCognitionSpontaneous2022. Furthermore, its role extends to translational applications, such as improving neural prosthetics and developing closed-loop BCIs, where spike sorting is critical for achieving precise neural decoding and control park128ChannelFPGABasedRealTime2017; hanLiveDemonstrationEfficient2024.
For a long time, spike sorting predominantly relied on heuristic statistics and machine learning approaches lewickiReviewMethodsSpike1998; quirogaUnsupervisedSpikeDetection2004; rossantSpikeSortingLarge2016; pachitariuFastAccurateSpike2016; chungFullyAutomatedApproach2017; hilgenUnsupervisedSpikeSorting2017; pachitariuSpikeSortingKilosort42024. While these methods have promoted neuroscience research, they exhibit several key limitations. First, their sorting results are sensitive to parameter settings and post-processing, which depend on the experimenter’s expertise and must be customized for each dataset. Moreover, these approaches lack a data-driven foundation, limiting their scalability and adaptability across diverse experimental settings, particularly in low signal-to-noise ratio (SNR) and high-variability scenarios, where their performance declines significantly. Recently, deep learning-based methods, such as YASS leeYASSAnotherSpike2017 and CEED vishnubhotlaRobustGeneralizableRepresentations2023, have attempted to improve spike sorting by adopting data-driven approaches to enhance detection and clustering accuracy. However, their generalizability and practical use are hindered by the limited training datasets.
@@ 2 Preliminaries
The goal of spike sorting (Figure 1) is to extract single-neuron spiking activity from extracellular recordings, where the activities of an unknown number of neurons are mixed together. Mathematically, spike sorting is a blind source separation problem buccinoSpikeSortingNew2022 and can be formally defined as follows: Let $\mathbf{V}\in\mathbb{R}^{T\times C}$ represent the extracellular recording (voltage) across $C$ electrode channels over $T$ time points. The goal is to infer a set of spike times $\{t_{k}\}_{k=1}^{K}$ and corresponding neuron labels $\{y_{k}\}_{k=1}^{K}$, where $y_{k}\in\{1,\dots,N\}$ represents the neuron identity, and $N$ is the total number of detected neurons.
@@ 3 Related Work
Early spike sorting pipelines typically used thresholds to detect spike events, followed by dimensionality reduction (e.g., PCA or $t$-SNE) and clustering FRS1901LIIIOL; maatenVisualizingDataUsing2008; veerabhadrappaCompatibilityEvaluationClustering2020. Although this pipeline was straightforward, it often suffered from noise susceptibility, inaccurate detection of low SNR events, and reliance on manual parameter tuning.
Deep learning approaches have been adopted to replace or enhance certain steps in traditional pipelines, aiming to improve robustness, automation, and adaptability to diverse experimental conditions. Autoencoders have been explored as a method for dimensionality reduction baldiAutoencodersUnsupervisedLearning2012; wuLearningSortFewshot2019; eomDeeplearnedSpikeRepresentations2021. YASS leeYASSAnotherSpike2017 employed a convolutional neural network for spike detection and waveform cleaning, thereby mitigating clustering errors caused by distorted waveforms. CEED vishnubhotlaRobustGeneralizableRepresentations2023 applied a contrastive learning framework to enforce invariances to amplitude fluctuations, noise, and channel subset changes in the extracted embeddings. However, these methods were limited by their reliance on restricted training datasets, potentially reducing their generalization across diverse experimental conditions.
@@ 4.1 Dataset Creation
To address the scarcity of labeled spike data, we generated a large-scale synthetic dataset using biophysically detailed neuron models from the Blue Brain Project (BBP) hayModelsNeocorticalLayer2011; markramReconstructionSimulationNeocortical2015; laquitaineSpikeSortingBiases2025, covering 206 neuron models from layers 1–6 of the juvenile rat somatosensory cortex. These models capture diverse morphologies and electrophysiological dynamics across 30 neuronal types.
@@
[…]
@@
We generated a large-scale dataset consisting of 8192 simulated recording trials, representing continuous neuronal activities from more than 40,000 individual neurons. Intracellular spike timestamps were recorded as ground-truth for extracellular spikes, providing precise annotations for model training and evaluation.
@@ 4.2 Spike Detection
We utilized a Transformer-based architecture vaswaniAttentionAllYou2017 for spike detection, trained on the simulated continuous signal dataset detailed in Sec. 4.1. For robustness and generalization, data augmentation was integrated into the training process, referring to Fig. 3. We employed a binary cross-entropy loss for the spike detection model:
@@ 4.3 Spike Identification
Contrastive Learning: We leveraged contrastive learning to learn representations that capture the relationships between spike waveforms (Fig. 4). Contrastive learning sought to map neural data into an embedding space where related examples were positioned nearby, while unrelated examples were placed further apart dorkenwaldMultilayeredMapsNeuropil2023; yuVivoCelltypeBrain2024. For each waveform $X_{i}$, we constructed a triplet by selecting $X_{i}^{+}$ with the same label and $X_{i}^{-}$ with a different label. The objective was schroffFaceNetUnifiedEmbedding2015:
@@
Denoising: The augmented waveform $X_{i}^{\prime}\in\mathbb{R}^{L\times C}$ then underwent a denoising step using Singular Value Decomposition (SVD) eckartApproximationOneMatrix1936; cunninghamDimensionalityReductionLargescale2014. The waveform was first reshaped and decomposed as $X_{i}^{\prime}=U\Sigma V^{\top}$, where $U$, $\Sigma$, and $V$ represented the singular vectors and singular values. The first $k$ components were retained to reconstruct the waveform.
Encoder: The denoised waveform $X_{i,\text{recon}}^{\prime}$ was fed into a Gated Recurrent Unit (GRU) encoder choLearningPhraseRepresentations2014 (Fig. 4), which processed the input sequentially. The final hidden state $h_{T}$ served as the learned representation $f(X_{i})$, encapsulating the spatiotemporal features of the waveform for subsequent clustering tasks.
Clustering: After obtaining the waveform representation, we applied dimensionality reduction using UMAP mcinnesUMAPUniformManifold2020 before clustering (Fig. 4). We tested two clustering approaches: a parametric algorithm, Gaussian Mixture Model (GMM) dempsterMaximumLikelihoodIncomplete1977, and a non-parametric one, Mean Shift (MS) chengMeanShiftMode1995.
@@ 5.1 Benchmarks
  • Real-world data-based synthetic datasets (with label): Hybrid (maglandSpikeForestReproducibleWebfacing2020), WaveClus (martinezRealisticSimulationExtracellular2009), IBL Neuropixels (vishnubhotlaRobustGeneralizableRepresentations2023). We evaluate both zero-shot (using only our simulated dataset) and fine-tuned (using also the dataset) performance of SimSort.
@@
Evaluation Metrics. To evaluate the performance of the overall spike sorting and spike detection tasks, we used Accuracy, Recall, and Precision as the evaluation metrics. Note that these metrics are specifically defined for the spike sorting context, where they measure the alignment between detected spikes and ground truth spike events, rather than their conventional definitions in classification tasks. For the spike identification task, we used the Adjusted Rand Index (ARI) hubert1985comparing to measure clustering accuracy relative to the ground truth.
Table 1: Spike sorting results on BBP L6 dataset. The results for other methods were obtained through SpikeInterface buccinoSpikeInterfaceUnifiedFramework2020. Values are as mean $\pm$ S.E.M.
@@ 5.2 Spike Sorting Results on simulated dataset
We first examined Simsort on a simulated test set (BBP L6). In Table 1, we evaluated SimSort alongside widely used algorithms, including KiloSort pachitariuFastAccurateSpike2016, KiloSort2, MountainSort4 chungFullyAutomatedApproach2017, and MountainSort5, implemented via the SpikeInterface pipeline buccinoSpikeInterfaceUnifiedFramework2020, on the simulated BBP L6 dataset. SimSort achieved the best performance across all metrics, demonstrating its robustness and strong zero-shot generalization ability under idealized, simulated conditions.
Table 2: Spike sorting results on Hybrid dataset. Data for other methods were sourced from SpikeForest maglandSpikeForestReproducibleWebfacing2020. Values are mean $\pm$ S.E.M. Best results are in bold. Note that SimSort did not use this dataset for training, indicating a zero-shot generalization.
Methods Hybrid-static (SNR>>3, 9 recordings) Hybrid-drift (SNR>>3, 9 recordings)
@@
[…]
Table 3: Spike sorting results on WaveClus dataset. The results for other methods were obtained through SpikeInterface buccinoSpikeInterfaceUnifiedFramework2020. Values are mean $\pm$ S.E.M. Note that SimSort did not use this dataset for training, indicating a zero-shot generalization.
@@ 5.3 Spiking Sorting results on real-world data-based synthetic datasets
Given SimSort’s strong performance on simulated data, we extended our evaluation to real-world data-based datasets to further test its zero-shot generalizability and robustness (RQ1). On the Hybrid dataset, we compared SimSort against several algorithms via SpikeForest, including HerdingSpikes2 hilgenUnsupervisedSpikeSorting2017, IronClust, JRClust, KiloSort pachitariuFastAccurateSpike2016, KiloSort2, MountainSort4 chungFullyAutomatedApproach2017, SpykingCircus, and Tridesclous. The results are summarized in Table 2 and Fig. S1. On this dataset, SimSort demonstrated the highest Accuracy and Precision in static recordings and achieved the best Accuracy and Recall under drift conditions, although it slightly trailed KiloSort in Precision.
@@ 5.5 Detailed Analysis
[…]
@@
After detecting spikes, our spike identification model extracted waveform representations in the latent space, enabling subsequent clustering. To effectively demonstrate the superiority of our identification model, we assembled a waveform dataset by gathering waveform data from each neuron in every hybrid recording. Using this dataset, we compared the 2D UMAP embedding of our model’s inferred representations with features extracted directly from waveforms using widely adopted dimensionality reduction techniques, including PCA FRS1901LIIIOL, $\boldsymbol{t}$-SNE maatenVisualizingDataUsing2008, and UMAP mcinnesUMAPUniformManifold2020. For evaluation, we applied a parametric clustering method, GMM dempsterMaximumLikelihoodIncomplete1977, to compute the Adjusted Rand Index (ARI), providing a quantitative measure of the reliability of the representations.
@@
Additionally, we compared SimSort with CEED vishnubhotlaRobustGeneralizableRepresentations2023, a recent framework for representation learning on extracellular data, using the author-provided model. Note that CEED is trained on datasets derived from IBL DanLab DY016 and DY009 Neuropixels recordings, which are more closely related to our test sets. We evaluated both SimSort and CEED on fifty test sets from the IBL Neuropixels dataset (specifically derived from the IBL CortexLab KS046 recording). As presented in Table S4, Fig. S2S3, our method, without training on the dataset, achieved performance comparable to CEED.
@@
Impact of Dimensionality Reduction and Clustering Methods: To evaluate the adaptability of SimSort, we conducted a study to assess the impact of different dimensionality reduction techniques (PCA, $t$-SNE, UMAP, and None) and clustering algorithms (GMM, KMeans macqueenMethodsClassificationAnalysis1967, Mean Shift chengMeanShiftMode1995, and Spectral Clustering ngSpectralClusteringAnalysis2001). Dimensionality reduction was applied to latent features extracted from waveforms, with an additional setup where clustering was performed directly on the raw latent features without dimensionality reduction (None). As illustrated in Fig. S4, the combination of UMAP and Mean Shift achieved the highest accuracy on both hybrid static and drift subsets. Across many other combinations, SimSort also maintained competitive performance, demonstrating its flexibility and robustness to various configurations.
@@ 6 Conclusion
This work was supported by AI for Science Foundation of Fudan University (FudanX24AI046 to Y.G.) and Microsoft Research.
Broader Impacts
Our work aims to advance the methodology of data processing in neuroscience experiments. It does not directly lead to any negative societal impact.